A random sample of 10 miniature Tootsie Rolls was taken from a bag. Each piece was weighed on a very accurate scale. The results in grams were

3.28
3.25
3.25
3.3
3.33
3.36
3.07
3.29
3.47
3.4


(a)Construct a 95 percent confidence interval for the true mean weight.

(b) What sample size would be necessary to estimate the true weight with an error of ± 0.02 grams with 95 percent confidence?

(c) What factors might cause variation in the weight of Tootsie Rolls during manufacture?


Answer by stanbon(75887)

A random sample of 10 miniature Tootsie Rolls was taken from a bag. Each piece was weighed on a very accurate scale. The results in grams were
3.28
3.25
3.25
3.3
3.33
3.36
3.07
3.29
3.47
3.4

(a)Construct a 95 percent confidence interval for the true mean weight.
x-bar: add them up and divide by 10 to get 3.3 and s = 0.1064
---
E ?: Using df=9 you get t= 2.262
E = 2.262*s/sqrt(n) = 2.262*0.1064/sqrt(10) = 0.07611
95% C.I.: 3.3-0.07611 < u < 3.3+0.0.07611
95% C.I.: 3.2239 < u < 3.3761
-------------------------------------------
(b) What sample size would be necessary to estimate the true weight with an error of ± 0.02 grams with 95 percent confidence?
sqrt(n) = z*0.1064/E
sqrt(n) = 1.96*0.064/0.02
n = 39.338
Rounding up: n = 40
===========================
(c) What factors might cause variation in the weight of Tootsie Rolls during manufacture?
I'll leave that up to your imagination.
==============================
Cheers,
Stan H.